# Can LLMs be Good Graph Judge for Knowledge Graph Construction?

### v3

@@ Abstract
Haoyu Huang1, Chong Chen2, Zeang Sheng3, Yang Li3, Wentao Zhang3 1The Chinese University of Hong Kong,2Huawei Cloud BU 3Peking University haoyuhuang@link.cuhk.edu.hk,chenchong55@huawei.com {shengzeang18,liyang.cs,wentao.zhang}@pku.edu.cn
@@ 1 Introduction
The transition from non-structured text to structured Knowledge Graphs (KGs) is a pivotal step in the evolution of data management and information retrieval systems. The task of automatic KG construction aims to develop a structured representation of knowledge from various data sources without the need for manual intervention. KGs usually serve as the backbone of numerous data science applications, including GraphRAG systems Edge et al. (2024) Peng et al. (2024) and recommendation systems Wang et al. (2019) Jiang et al. (2024). Exploring a way to construct high-quality KGs from unstructured data is crucial for different downstream applications based on KG Ge et al. (2021) Huang et al. (2024) Wei et al. (2024) Rabbani et al. (2023).
@@
Recently, Large Language Models (LLMs) have demonstrated significant generalization capabilities in various Natural Language Processing (NLP) tasks Pan et al. (2024) and KG related tasks, such as text generation Li et al. (2024), KG Completion (KGC) Yao et al. (2023) and Open Information Extraction (OpenIE) Angeli et al. (2015) Dagdelen et al. (2024). Consequently, there are many works that utilize LLMs to construct KGs from unstructured natural language documents. The incorporation of LLMs can address the issue of generalization in open-domain applications Carta et al. (2023a). With its robust zero-shot generation capability, there is no need for us to gather a large volume of annotated data for tasks such as named entity recognition (NER), entity extraction, or relation extraction.
@@
Although recent LLM-based methods Mo et al. (2025) Han et al. (2023) Lairgi et al. (2024) have gained some success in the KG construction task, we find that they may still face three challenges:
@@
(1) Noise Information. Real-world documents are not only voluminous but also rife with noise, which poses a significant challenge for LLMs extracting valuable structured information. The sheer volume of data can lead to the extraction of excessive and irrelevant information, overshadowing the critical insights that LLMs are meant to uncover Liu et al. (2024b) Shi et al. (2023a). For example, as shown in Figure 1, the triple <Protein X, is on, a 50% Discount>is incorrectly constructed due to the irrelevant advertisement with red lines in the document, which is the noise information that makes the LLM incorrectly believed that the Protein X is on sale with discounts.
@@
(2) Domain-Specific Knowledge. Naive LLMs often generate inaccurate triples with domain-specific documents, which require a deep understanding of specialized terminology and context Zhong et al. (2023) Zhu et al. (2024). And this kind of error is hard to be observed by naive LLMs. For example, in Figure 1, the triple <Protein X, cured, Patient>is inaccurately extracted due to a lack of medical domain-specific knowledge. While the document marks a reference with blue lines, note that the original text only suggests a link between Protein X and better patient results, not that it can cure patients in medical fields.
@@
(3) Hallucinations of LLMs. When LLMs are directly used to build KGs, they are prone to generating false or distorted information, which is a phenomenon called hallucinations Zhang et al. (2023) Ji et al. (2023). This can lead to the incorporation of inaccurate or fabricated facts into the KG, undermining the reliability of the KG. For example, as shown in Figure 1, the triple marked in green <AI Technology, replaces, Chemotherapy>is incorrectly generated without any reference in the original document, even in the entity-related text highlighted with a green line.
@@
To this end, we propose a new method called GraphJudge, which utilizes a fine-tuned open source LLM (e.g., LLaMA-2 Touvron et al. (2023)) as an expert to judge the correctness of the triples generated by another closed-source LLM (e.g., GPT-4o-mini). To address the first challenge, we introduce the Entity-Centric Text Denoising (ECTD) module. We clean up the original documents by eliminating redundant words and irrelevant information not pertinent to the entities identified by the LLM. This module also leverages the robust zero-shot generation capabilities of LLMs to ensure the recall of a sufficient number of triple candidates Wei et al. (2023) Carta et al. (2023a). To overcome the second challenge, we suggest the module of Knowledge Aware Supervised Fine-Tuning (KASFT). We introduce the graph judgement task from the triple classification task. To verify the accuracy of the triples generated by the closed-source LLM, we conduct supervised fine-tuning (SFT) on an open-source LLM, which can make it achieve over 90% accuracy on graph judgement tasks with strong generalization abilities. To settle the third challenge, the Graph Judgement (GJ) module is introduced. We utilize the fine-tuned open-source LLM to conduct judgement on the generated triples in the first module and filter out the wrong items to finally improve the quality of generated KGs.
@@ 2 Related Work
In this section, we will introduce recent LLM-based OpenIE and KG construction methods. Some workAgrawal et al. (2022) Wei et al. (2023) has demonstrated that LLMs have remarkable zero-shot and few-shot information extraction abilities. However, they face difficulties when it comes to more intricate tasks such as relation extraction and event extraction Carta et al. (2023a). To address that, Kumar et al. Kumar et al. (2020) propose a unified approach to construct KGs from unprocessed text. They initially fine-tuned a pre-trained language model (PLM) for NER. Subsequently, they introduced a ”2-model BERT” architecture to extract relations. GPT-RE Wan et al. (2023) introduces the in-context learning method and task-aware representations in demonstration retrieval and aims to enhance the connections between examples and triples. PiVe Han et al. (2023) designs a paradigm that fine-tuning a PLM as the verifier to predict the missing triples. With iterative verifications, the graph-based generative capability of LLMs can be improved. VicunaNER Ji (2023) utilizes the open-source LLM Vicuna to do zero-shot or few-shot NER. Similarly, it also performs recognition to identify entities that were not recognized in the previous phase. Carta et al. Carta et al. (2023a) develops an iterative LLM prompting-based pipeline to generate KGs without requiring predefined sets or external ontologies. iText2KG Lairgi et al. (2024) proposes a zero-shot method to construct consistent KGs from documents with LLMs. It restructures the unprocessed documents using a preset template and identifies distinct entities and connections in a semantic manner. SAC-KG Chen et al. (2024) exploits LLMs as skilled automatic constructors for domain KGs and employs a naive LLM to predict the correctness of constructed triples. KGGen Mo et al. (2025) clusters related entities to reduce sparsity in the KGs constructed by LLMs.
@@ 4.2.1 Text denoising and entity extraction
In phrase 1, we consider that a substantial portion of real-world documents retrieved from information retrieval systems are consist of considerable noise information. And that may influence the quality of relations extracted by LLM Shi et al. (2023b) Liu et al. (2024c). So we design an iterative denoising method to remove messy information from the original text.
Specifically, we extract entities from the original document using LLM. Subsequently, we input these entities and the original document into LLM to generate the denoised document. In this way, we can achieve two goals: (1) The noise information that is not related to the topic of the document can be removed. (2) The content of the documents can be reorganized in an entity-centric way, which is friendly to the triple extraction in the next phrase. Finally, for each raw document we will get the extracted entity set and the denoised document . Note that important information can be well preserved in as verified in Appendix F.
@@ 4.3 Knowledge Aware Supervised Fine-Tuning
In this module, inspired by KG-LLaMA Yao et al. (2023), we propose the method of treating triples in the draft KG as textual sequences and model graph judgement task as a sequence-to-sequence problem. We construct instruction data from the training set and fine-tune an open-source LLM to achieve the goal of both excelling at checking facts from documents with the triple structure and acknowledgment of domain-specific knowledge. The LLM can also learn how to verify the consistency between the document and the extracted triples. Checking facts with the triple structure refers to the general structure of triples is often analogous to a grammatical subject, predicate, and object or a subject with a relational attribute. LLMs are anticipated to have the ability to identify their correctness from the give documents. Domain-specific knowledge refers to the knowledge in the documents could be a new domain Zhong et al. (2023), which is typically not part of pre-training data of LLMs. By employing SFT, the domain-specific knowledge from the documents can be incorporated into the LLM, thus enhances its graph judgment performance. And only if LLMs are fine-tuned as graph judges, these types of knowledge can be well learned, as justified in Figure 5 and Appendix G.
@@
Furthermore, we transfer the triples in to natural language sentences to construct the instruction data with paired documents as contexts following the prompt templates shown in Appendix H. The triple sentences either represent a real fact or a fake fact. Then let the LLM make judgements with these instructions. Mathematically, with tokenized sentences transferred from triples and paired documents , and tokenized instruction , for a sequence of length , we compute the probability of generating the target output as follows:
@@ 4.4 Graph Judgement
The KGs created in the first module are preliminary and that is also why we call that draft KGs. In this module, we will verify the triples in these draft KGs using our fine-tuned LLM in the second module.
@@ 5.1.1 Dataset
In our study, we conduct experiments on two general datasets (REBEL-Sub Huguet Cabot and Navigli (2021) and GenWiki Jin et al. (2020)) and one domain-specific dataset (SCIERC Luan et al. (2018)) with golden ground truth KGs. We demonstrate the detailed information and statistics of each dataset in Appendix A. For each dataset we randomly select a sample of 2000 data points from the training data for validation purposes during the fine-tuning of the LLM.
@@ 5.1.2 Baselines
In our performance comparison, we consider two methods for comprehensive evaluation: GPT-4o-mini: We conduct experiments on GPT-4o with one-shot learning method. The instructions we have developed are identical to those outlined in our method. GPT-4o Hurst et al. (2024): The same settings as GPT-4o-mini. PiVe Han et al. (2023): We follow the default parameter settings of PiVe. We use the largest verifier module in PiVe, Flan-T5-XXL Chung et al. (2024). We employ the LoRA adapter checkpoint111https://huggingface.co/Jiuzhouh/flan-t5-xxl-lora-verifier, which has been well trained. And the LLM we use in this model is GPT-4o-mini. We implement an iterative prompting approach with three rounds, which represents the optimal number of iteration rounds as outlined in their study. KGGen Mo et al. (2025): We also employ GPT-4o-mini as the LLM used in this baseline.
@@ 5.1.3 Implementation Details
| GraphJudge | 0.6847 | 0.8775 | 0.7283 | 0.4898 | 0.6273 | 0.5216 | 0.4321 | 0.5591 | 0.4603 |
@@
Supervised Fine-Tuning: We employ LLaMA-2-7B as the base model to carry out SFT with LoRA Hu et al. (2021). The instructions are constructed with the documents, query sentences, and the triple sentences. We perform SFT on autoregression generation tasks, which is a common approach to fine-tune LLMs Black et al. (2022). The expected responses (labels) are either ”Yes, that is true.” or ”No, that is not true.”. Training settings are illustrated in Appendix C. The training was done using a single L20 GPU with 48GB of RAM.
@@ 5.2 Overall Performance Comparison (RQ1)
GraphJudge’s superior performance. GraphJudge outperforms other baselines in most of the cases. The superiority of GraphJudge’s F1 scores (marked with gray color) demonstrates that, while maintaining a reasonable level of recall for triples, it also achieves improvement in accuracy. For example, as the results marked with red color show, although PiVe exhibits stronger recall ability, it overlooks triple accuracy. KGGen excels in accuracy but fails at recall. In contrast, GraphJudge leverages the ECTD module based on a closed-source LLM to ensure recall ability, while the KASFT and GJ modules with a fine-tuned open-source LLM guarantee accuracy, enabling its F1 score to surpass those of other baseline models. We can also observe that GraphJudge excels not only with domain-specific documents, but also demonstrates superior performance with general documents.
@@ 5.3 Module Ablation Study (RQ2)
Effect of Entity-Centric Text Denoising. We investigate the benefit of introducing entity-centric denoising paradigm using the variant ”w/o ECTD”, where we do not conduct document denoising and directly extract entities and relations from original documents. The results show that our full model performs significantly better than this ablated version. It suggests that ECTD module can avoid LLMs extract wrong structured information from irrelevance or not well-formatted corpus.
@@
Furthermore, to showcase the noise reduction capability of ECTD, we visualize the semantic correlation of the triples in a known KG with the denoised and original document, respectively. As shown in Figure 4, deeper color in the heat maps suggests a stronger relevance. The refined document exhibits greater relevance to the triples, demonstrating the effectiveness of ECTD. Implementation details are described in Appendix E.
@@
Effect of Knowledge Aware Supervised Fine-Tuning. We conduct graph judgement on the triples without fine-tuning the open-source LLM, which is denoted as ”w/o KASFT”. The result in Table 2 indicates that without SFT, the naive LLM has weak graph judgement abilities. And with a fine-tuned LLM as a graph judge, the performance can be improved a lot. Because KASFT enables the LLM to acquire both fact-checking capabilities and domain-specific knowledge within the triples in our instruction training data.
@@
Furthermore, we apply negative sampling to construct instructions on the test set like what we did on the training set. We randomly select 500 samples and perform graph judgement to compare the capabilities of different models. As shown in Figure 5 and Appendix G, both fine-tuned small models like BERT and naive powerful LLMs like GPT-4o show poor performance on the graph judgement task even with documents as contexts. However, GraphJudge can achieve over 90% judgement accuracy on REBEL-Sub and GenWiki, which demonstrates the KASFT module can indeed enhance the effectiveness of LLMs as a graph judge.
@@
Effect of Graph Judgement. We compare the performances of our full model and the model without GJ module denoted as ”w/o GJ”. The result suggests that GJ module plays a very important role in GraphJudge. It can significantly enhance the quality of KGs generated by the closed-source LLM and reduce the effects of the inaccuracies or hallucinations that may arise from LLMs. The closed-source LLM excels in zero-shot generation, boosting recall but suffering accuracy due to hallucinations or knowledge inadequacy. The GJ module relieves this by filtering inaccurate triples, enhancing the quality of constructed KGs.
| Method | GenWiki @ REBEL-Sub | ||
| G-BS-F1↑ | G-BL-F1↑ | G-RO-F1↑ | |
| GPT-4o | 0.4163 | 0.3158 | 0.2966 |
| PiVe | 0.4090 | 0.3010 | 0.2823 |
| KGGen | 0.4995 | 0.3146 | 0.2719 |
| GraphJudge | 0.5814 | 0.4055 | 0.3649 |

### v4

@@ Abstract
Haoyu Huang1, Chong Chen2, Zeang Sheng3, Yang Li3, Wentao Zhang3 1Hong Kong University of Science and Technology,2Huawei Cloud BU 3Peking University hhuangcp@connect.ust.hk,chenchong55@huawei.com {shengzeang18,liyang.cs,wentao.zhang}@pku.edu.cn
@@ 1 Introduction
The transition from non-structured text to structured Knowledge Graphs (KGs) is a pivotal step in the evolution of data management and information retrieval systems. The task of automatic KG construction aims to develop a structured representation of knowledge from various data sources without the need for manual intervention. KGs usually serve as the backbone of numerous data science applications, including GraphRAG systems Edge et al. (2024); Peng et al. (2024); Huang et al. (2025) and recommendation systems Wang et al. (2019); Jiang et al. (2024); Chen et al. (2025). Exploring a way to construct high-quality KGs from unstructured data is crucial for different downstream applications based on KG Ge et al. (2021); Huang et al. (2024); Wei et al. (2024); Rabbani et al. (2023).
@@
Recently, Large Language Models (LLMs) have demonstrated significant generalization capabilities in various Natural Language Processing (NLP) tasks Pan et al. (2024) and KG related tasks, such as text generation Li et al. (2024), KG Completion (KGC) Yao et al. (2023) and Open Information Extraction (OpenIE) Angeli et al. (2015); Dagdelen et al. (2024). Consequently, there are many works that utilize LLMs to construct KGs from unstructured natural language documents. The incorporation of LLMs can address the issue of generalization in open-domain applications Carta et al. (2023a). With its robust zero-shot generation capability, there is no need for us to gather a large volume of annotated data for tasks such as named entity recognition (NER), entity extraction, or relation extraction.
@@
Although recent LLM-based methods Mo et al. (2025); Han et al. (2023); Lairgi et al. (2024) have gained some success in the KG construction task, we find that they may still face three challenges:
@@
(1) Noise Information. Real-world documents are not only voluminous but also rife with noise, which poses a significant challenge for LLMs extracting valuable structured information. The sheer volume of data can lead to the extraction of excessive and irrelevant information, overshadowing the critical insights that LLMs are meant to uncover Liu et al. (2024b); Shi et al. (2023a). For example, as shown in Figure 1, the triple <Protein X, is on, a 50% Discount>is incorrectly constructed due to the irrelevant advertisement with red lines in the document, which is the noise information that makes the LLM incorrectly believed that the Protein X is on sale with discounts.
@@
(2) Domain-Specific Knowledge. Naive LLMs often generate inaccurate triples with domain-specific documents, which require a deep understanding of specialized terminology and context Zhong et al. (2023); Zhu et al. (2024). And this kind of error is hard to be observed by naive LLMs. For example, in Figure 1, the triple <Protein X, cured, Patient>is inaccurately extracted due to a lack of medical domain-specific knowledge. While the document marks a reference with blue lines, note that the original text only suggests a link between Protein X and better patient results, not that it can cure patients in medical fields.
@@
(3) Hallucinations of LLMs. When LLMs are directly used to build KGs, they are prone to generating false or distorted information, which is a phenomenon called hallucinations Zhang et al. (2023); Ji et al. (2023). This can lead to the incorporation of inaccurate or fabricated facts into the KG, undermining the reliability of the KG. For example, as shown in Figure 1, the triple marked in green <AI Technology, replaces, Chemotherapy>is incorrectly generated without any reference in the original document, even in the entity-related text highlighted with a green line.
@@
To this end, we propose a new method called GraphJudge, which utilizes a fine-tuned open source LLM (e.g., LLaMA-2 Touvron et al. (2023)) as an expert to judge the correctness of the triples generated by another closed-source LLM (e.g., GPT-4o-mini). To address the first challenge, we introduce the Entity-Centric Text Denoising (ECTD) module. We clean up the original documents by eliminating redundant words and irrelevant information not pertinent to the entities identified by the LLM. This module also leverages the robust zero-shot generation capabilities of LLMs to ensure the recall of a sufficient number of triple candidates Wei et al. (2023); Carta et al. (2023a). To overcome the second challenge, we suggest the module of Knowledge Aware Supervised Fine-Tuning (KASFT). We introduce the graph judgement task from the triple classification task. To verify the accuracy of the triples generated by the closed-source LLM, we conduct supervised fine-tuning (SFT) on an open-source LLM, which can make it achieve over 90% accuracy on graph judgement tasks with strong generalization abilities. To settle the third challenge, the Graph Judgement (GJ) module is introduced. We utilize the fine-tuned open-source LLM to conduct judgement on the generated triples in the first module and filter out the wrong items to finally improve the quality of generated KGs.
@@ 2 Related Work
In this section, we will introduce recent LLM-based OpenIE and KG construction methods. Some workAgrawal et al. (2022); Wei et al. (2023) has demonstrated that LLMs have remarkable zero-shot and few-shot information extraction abilities. However, they face difficulties when it comes to more intricate tasks such as relation extraction and event extraction Carta et al. (2023a). To address that, Kumar et al. Kumar et al. (2020) propose a unified approach to construct KGs from unprocessed text. They initially fine-tuned a pre-trained language model (PLM) for NER. Subsequently, they introduced a ‘2-model BERT’ architecture to extract relations. GPT-RE Wan et al. (2023) introduces the in-context learning method and task-aware representations in demonstration retrieval and aims to enhance the connections between examples and triples. PiVe Han et al. (2023) designs a paradigm that fine-tuning a PLM as the verifier to predict the missing triples. With iterative verifications, the graph-based generative capability of LLMs can be improved. VicunaNER Ji (2023) utilizes the open-source LLM Vicuna to do zero-shot or few-shot NER. Similarly, it also performs recognition to identify entities that were not recognized in the previous phase. Carta et al. Carta et al. (2023a) develops an iterative LLM prompting-based pipeline to generate KGs without requiring predefined sets or external ontologies. iText2KG Lairgi et al. (2024) proposes a zero-shot method to construct consistent KGs from documents with LLMs. It restructures the unprocessed documents using a preset template and identifies distinct entities and connections in a semantic manner. SAC-KG Chen et al. (2024) exploits LLMs as skilled automatic constructors for domain KGs and employs a naive LLM to predict the correctness of constructed triples. KGGen Mo et al. (2025) clusters related entities to reduce sparsity in the KGs constructed by LLMs.
@@ 4.2.1 Text denoising and entity extraction
In phrase 1, we consider that a substantial portion of real-world documents retrieved from information retrieval systems are consist of considerable noise information. And that may influence the quality of relations extracted by LLM Shi et al. (2023b); Liu et al. (2024c). So we design an iterative denoising method to remove messy information from the original text.
Specifically, we extract entities from the original document using LLM. And as verified in Appendix E, the entities extracted by closed-source LLMs have a high coverage rate and provide a good foundation for the following denoising, relation extraction and triple filtering processes. Subsequently, we input these entities and the original document into LLM to generate the denoised document. In this way, we can achieve two goals: (1) The noise information that is not related to the topic of the document can be removed. (2) The content of the documents can be reorganized in an entity-centric way, which is friendly to the triple extraction in the next phrase. Finally, for each raw document we will get the extracted entity set and the denoised document . Note that important information can be well preserved in as verified in Appendix G.
@@ 4.3 Knowledge Aware Supervised Fine-Tuning
In this module, inspired by KG-LLaMA Yao et al. (2023), we propose the method of treating triples in the draft KG as textual sequences and model graph judgement task as a sequence-to-sequence problem. We construct instruction data from the training set and fine-tune an open-source LLM to achieve the goal of both excelling at checking facts from documents with the triple structure and acknowledgment of domain-specific knowledge. The LLM can also learn how to verify the consistency between the document and the extracted triples. Checking facts with the triple structure refers to the general structure of triples is often analogous to a grammatical subject, predicate, and object or a subject with a relational attribute. LLMs are anticipated to have the ability to identify their correctness from the give documents. Domain-specific knowledge refers to the knowledge in the documents could be a new domain Zhong et al. (2023), which is typically not part of pre-training data of LLMs. By employing SFT, the domain-specific knowledge from the documents can be incorporated into the LLM, thus enhances its graph judgment performance. And only if LLMs are fine-tuned as graph judges, these types of knowledge can be well learned, as justified in Figure 5 and Appendix H.
@@
Furthermore, we transfer the triples in to natural language sentences to construct the instruction data with paired documents as contexts following the prompt templates shown in Appendix I. The triple sentences either represent a real fact or a fake fact. Then let the LLM make judgements with these instructions. Mathematically, with tokenized sentences transferred from triples and paired documents , and tokenized instruction , for a sequence of length , we compute the probability of generating the target output as follows:
@@ 4.4 Graph Judgement
The KGs created in the first module are preliminary and that is also why we call that draft KGs. In this module, we will judge the factual correctness of the triples in these draft KGs using our fine-tuned LLM in the second module and filter out the incorrect triples.
@@ 5.1.1 Dataset
In our study, we conduct experiments on two general datasets (REBEL-Sub Huguet Cabot and Navigli (2021) and GenWiki Jin et al. (2020)) and two domain-specific datasets (SCIERC Luan et al. (2018) and the Windows-centric subset of Re-DocRED Tan et al. (2022) used by Sun et al. (2025)) with golden ground truth KGs. We demonstrate the detailed information and statistics of each dataset in Appendix A. For each dataset we randomly select a sample of 2000 data points from the training data for validation purposes during the fine-tuning of the LLM.
@@ 5.1.2 Baselines
In our performance comparison, we consider six baselines for comprehensive evaluation: GPT-4o-mini: We conduct experiments on GPT-4o with one-shot learning method. The instructions we have developed are identical to those outlined in our method. GPT-4o Hurst et al. (2024): The same settings as GPT-4o-mini. RAKG Zhang et al. (2025), iText2KG Lairgi et al. (2024), and KGGen Mo et al. (2025): We follow the default settings of them and their official implementations with GPT-4o-mini as the LLM. PiVe Han et al. (2023): We follow the default parameter settings of PiVe. We use the largest verifier module in PiVe, Flan-T5-XXL Chung et al. (2024). We employ the LoRA adapter checkpoint111https://huggingface.co/Jiuzhouh/flan-t5-xxl-lora-verifier, which has been well trained. And the LLM we use in this model is GPT-4o-mini. We implement an iterative prompting approach with three rounds, which represents the optimal number of iteration rounds as outlined in their study.
@@ 5.1.3 Implementation Details
| GraphJudge | 0.6847 | 0.8775 | 0.7283 | 0.4898 | 0.6273 | 0.5216 | 0.4321 | 0.5591 | 0.4603 | |
| Re-DocRED | GPT-4o-mini | 0.8036 | 0.7460 | 0.6807 | 0.4840 | 0.4659 | 0.4254 | 0.3337 | 0.3338 | 0.2938 |
| GPT-4o | 0.7508 | 0.7588 | 0.6864 | 0.4438 | 0.4682 | 0.4066 | 0.3076 | 0.3383 | 0.2833 | |
| RAKG | 0.3121 | 0.9243 | 0.4422 | 0.2144 | 0.6377 | 0.3044 | 0.1613 | 0.4922 | 0.2306 | |
| iText2KG | 0.7957 | 0.5519 | 0.5502 | 0.4930 | 0.3741 | 0.3522 | 0.3704 | 0.2958 | 0.2678 | |
| PiVe | 0.7963 | 0.6769 | 0.6289 | 0.4545 | 0.4193 | 0.3729 | 0.2848 | 0.2955 | 0.2506 | |
| KGGen | 0.8422 | 0.3914 | 0.4372 | 0.5284 | 0.2594 | 0.2768 | 0.3662 | 0.1937 | 0.1959 | |
| GraphJudge | 0.7801 | 0.7579 | 0.7051 | 0.4776 | 0.4830 | 0.4322 | 0.3350 | 0.3532 | 0.3048 |
@@
Supervised Fine-Tuning: We employ LLaMA-2-7B as the base model to carry out SFT with LoRA Hu et al. (2021). The instructions are constructed with the documents, query sentences, and the triple sentences. We perform SFT on autoregression generation tasks, which is a common approach to fine-tune LLMs Black et al. (2022). The expected responses (labels) are either ‘Yes, that is true.’ or ‘No, that is not true.’. Training settings are illustrated in Appendix C. The training was done using a single L20 GPU with 48GB of RAM.
@@ 5.2 Overall Performance Comparison (RQ1)
GraphJudge’s superior performance. GraphJudge outperforms other baselines in most of the cases. The superiority of GraphJudge’s F1 scores (marked with gray color) demonstrates that, while maintaining a reasonable level of recall for triples, it also achieves improvement in accuracy. For example, as the results marked with red color show, although RAKG and PiVe exhibit stronger recall ability, they overlook triple accuracy. KGGen excels in accuracy but fails at recall. In contrast, GraphJudge leverages the ECTD module based on a closed-source LLM to ensure recall ability, while the KASFT and GJ modules with a fine-tuned open-source LLM guarantee accuracy, enabling its F1 score to surpass those of other baseline models. We can also observe that GraphJudge excels not only with domain-specific documents, but also demonstrates superior performance with general documents.
@@ 5.3 Module Ablation Study (RQ2)
Effect of Entity-Centric Text Denoising. We investigate the benefit of introducing entity-centric denoising paradigm using the variant ‘w/o ECTD’, where we do not conduct document denoising and directly extract entities and relations from original documents. The results show that our full model performs significantly better than this ablated version. It suggests that ECTD module can avoid LLMs extract wrong structured information from irrelevance or not well-formatted corpus.
@@
Furthermore, to showcase the noise reduction capability of ECTD, we visualize the semantic correlation of the triples in a known KG with the denoised and original document, respectively. As shown in Figure 4, deeper color in the heat maps suggests a stronger relevance. The refined document exhibits greater relevance to the triples, demonstrating the effectiveness of ECTD. Implementation details are described in Appendix F.
@@
Effect of Knowledge Aware Supervised Fine-Tuning. We conduct graph judgement on the triples without fine-tuning the open-source LLM, which is denoted as ‘w/o KASFT’. The result in Table 2 indicates that without SFT, the naive LLM has weak graph judgement abilities. And with a fine-tuned LLM as a graph judge, the performance can be improved a lot. Because KASFT enables the LLM to acquire both fact-checking capabilities and domain-specific knowledge within the triples in our instruction training data.
@@
Furthermore, we apply negative sampling to construct instructions on the test set like what we did on the training set. We randomly select 500 samples and perform graph judgement to compare the capabilities of different models. As shown in Figure 5 and Appendix H, both fine-tuned small models like BERT and naive powerful LLMs like GPT-4o show poor performance on the graph judgement task even with documents as contexts. However, GraphJudge can achieve over 90% judgement accuracy on REBEL-Sub and GenWiki, which demonstrates the KASFT module can indeed enhance the effectiveness of LLMs as a graph judge.
| Method | GenWiki @ REBEL-Sub | REBEL-Sub @ GenWiki | REBEL-Sub @ SCIERC | ||||||
| G-BS-F1↑ | G-BL-F1↑ | G-RO-F1↑ | G-BS-F1↑ | G-BL-F1↑ | G-RO-F1↑ | G-BS-F1↑ | G-BL-F1↑ | G-RO-F1↑ | |
| GPT-4o | 0.4163 | 0.3158 | 0.2966 | 0.8428 | 0.6774 | 0.6028 | 0.7035 | 0.5032 | 0.4425 |
| PiVe | 0.4090 | 0.3010 | 0.2823 | 0.8230 | 0.6503 | 0.5817 | 0.6725 | 0.4924 | 0.4385 |
| KGGen | 0.4995 | 0.3146 | 0.2719 | 0.8169 | 0.5542 | 0.4641 | 0.6635 | 0.4725 | 0.4211 |
| GraphJudge | 0.5814 | 0.4055 | 0.3649 | 0.8587 | 0.6792 | 0.5911 | 0.7431 | 0.5156 | 0.4572 |
@@
Effect of Graph Judgement. We compare the performances of our full model and the model without GJ module denoted as ‘w/o GJ’. The result suggests that GJ module plays a very important role in GraphJudge. It can significantly enhance the quality of KGs generated by the closed-source LLM and reduce the effects of the inaccuracies or hallucinations that may arise from LLMs. The closed-source LLM excels in zero-shot generation, boosting recall but suffering accuracy due to hallucinations or knowledge inadequacy. The GJ module relieves this by filtering inaccurate triples, enhancing the quality of constructed KGs.